C. RELATION.

Given the following set:

1. X = {1, 2, 3, 4, 5} defined by the rule (x, y) ∈ R if x + y ≤ 6


a. List the elements of R

b. Find the domain of R

c. Find the range of R

d. Draw the digraph

e. Properties of the Relation 


1
Expert's answer
2021-03-15T12:20:36-0400

Let's list the elements of R:

"R = \{ (1,1),\,(1,2),\,(1,3),\,(1,4),\,(1,5),\,(2,1),\,(2,2),\,(2,3),\,(2,4),\,(3,1),\,(3,2),\,(3,3),\,(4,1),\,(4,2),\,(5,1)\}"

Find the domain  of the relation:

"D(R) = \{ x|(x,y) \in R\} = \{ 1,2,3,4,5\} = X"

Find the range  of the relation:

"E(R) = \{ y|(x,y) \in R\} = \{ 1,2,3,4,5\} = X"

Draw the digraph:



Find properties of the Relation:

"(5,5) \notin R" so R is not reflexive

"(1,1) \in R" so R is not irreflexive

If "x + y \le 6" then "y+x \le 6" so "\left( {x,y} \right) \in R \Leftrightarrow \left( {y,x} \right) \in R" - R is symmetric relation.

"(5,1) \in R,\,(1,4) \in R" , but "(5,4) \notin R" so R is not transitive relation.